OpenStudy (mitu12):

find the greatest common factor between these two terms 39v and 30uv. i need to know how to do these type of problems.

OpenStudy (anonymous):

"common" means "the same" in 39v, and 30uv, first glance gives us the "common" letter is ???

OpenStudy (mitu12):

v

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

now, number: if I write 39 = 3*13 and 30 = 3*10 does it make sense?

OpenStudy (mitu12):

yes

OpenStudy (anonymous):

so, the "common" number among them is??

OpenStudy (mitu12):

3

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Now, combine: 39v = 3*13*v 30uv = 3*10*u*v the "common" is??

OpenStudy (mitu12):

3 and v

OpenStudy (anonymous):

3v, not 3 and v

OpenStudy (anonymous):

got it?

OpenStudy (mitu12):

yes

OpenStudy (anonymous):

done :)

OpenStudy (mitu12):

thank you could you give me a an example problem please

OpenStudy (anonymous):

Find the greatest common factor between these terms: 4xy and 2y

OpenStudy (mitu12):

2y

OpenStudy (anonymous):

yup Now, find the greatest common factor among 3abc , bc, 4ac

OpenStudy (mitu12):

c

OpenStudy (mitu12):

@OOOPS

OpenStudy (anonymous):

Perfect!!

OpenStudy (mitu12):

thank you so much

OpenStudy (anonymous):

:)